Find the equation of the cone with the vertex at (1,−1,2) and the base curve as
(z+1)
2 = x+2, y = 3.
1
Expert's answer
2020-03-09T13:48:46-0400

Solution:

The vertex "V(1,-1,2)". Let a, b, c be the direction ratios of generator of the cone.

Then the equations of generator are

"\frac{x-1}{a}=\frac{y+1}{b}=\frac{z-2}{c}=t" (say) (I)

The coordinates of any points on the generator are "(1+at, -1+bt, 2+ct)".

For some "t \in \Reals, (1+at, -1+bt, 2+ct)" lies on the guiding curve.

Therefore "((2+ct)+1)^2=(1+at)+2" and "-1+bt=3". Thus, "t=\frac{4}{b}".

From this we get

"((2+c\cdot\frac{4}{b})+1)^2=(1+a\cdot\frac{4}{b})+2"

From (I) we obtain:

"(3+4\cdot\frac{z-2}{y+1})^2=3+4\cdot\frac{x-1}{y+1}" .

After simplification, the required equation of the cone is

"6y^2+24yz-4xy-32y+16z^2+26-40z-4x=0".

Answer:

"6y^2+24yz-4xy-32y+16z^2+26-40z-4x=0".